David and Ruby are engaged and plan to get married. During 2017, David is a full-time student and earns $9,400 from a part-time job. With this income, student loans, savings, and nontaxable scholarships, he is self-supporting. For the year, Ruby is employed and reports $74,600 in wages. The personal exemption amount for 2017 is $4,050.

Click here to access the standard deduction table to use. Click here to access the Tax Rate Schedules. If an amount is zero, enter, "0". Do not round your intermediate computations. Round your final answer to nearest whole dollar.

a. Compute the following:

David
Filing Single
Ruby
Filing Single
Gross income and AGI $ $
Standard deduction
Personal exemption
Taxable income $ $
Income tax $ $

Feedback

If married individuals elect to file separate returns, each reports only his or her own income, exemptions, deductions, and credits, and each must use the Tax Rate Schedule applicable to married taxpayers filing separately. It is generally advantageous for married individuals to file a joint return, because the combined amount of tax is lower.

b. Assume that David and Ruby get married in 2017 and file a joint return. What is their taxable income and income tax? Round your final answer to nearest whole dollar.

Married
Filing Jointly
Gross income $
Standard deduction
Personal exemptions
Taxable income $
Income tax $

c. How much Federal income tax can David and Ruby save if they get married in 2017 and file a joint return?
$
Answer #1

Answer (a):

Gross income and AGI Standard deduction Personal exemption Taxable income Income tax David Filing Single $9,400 $6,350 $4,050

For Ruby tax calculation is as follows:

Tax = 5226.25 + (64200 - 37950)* 25%

= $11788.25

Answer (b):

Gross income Standard deduction Personal exemptions Taxable income Income tax Married Filing Jointly $84,000 $12,700 $8,100 $

Working:

Joint gross income = 9400 + 74600 = $84,000

MFJ:

Tax for 2017 = 1865 + (63200 - 18650) * 15% = $8547.50

Answer (c):

Federal income tax David and Ruby can save if they get married in 2017 and file a joint return = (0 + 11789) - 8548

= $3,241

Federal income tax David and Ruby can save if they get married in 2017 and file a joint return = $3,241